1) The slope and intercept of an Arrhenius plot made for the first-order decomposition reaction are -2864.3 K and 23.279 at 303K

2) The value of Keq for the following reaction is 0.340: A (g) +B (g)=C(g) +D (g) The value of Keq at the same temperature for t

3) 2NO(g) +Cl2 (g) → NOCI (g) Experiment concentration of NO (M) 0.060 0.015 0.015 concentration of Cl2(M) Rate (M/s) 0.024 3.40
Answer #1

Arrhenius equation is K = ė EART taking log both side. loge = log A - Ea a Intercept (c) = logA = 23:279 slope (m) = - ca. =A & B ²c +D. Keq = .34. CAD Key - 3 Ž A&B 30+ 3D 23 A+ 3B keq = hE. Keq=25:42het rate of reaction. 8= k. (Noy a lazy B. order of reaction. = x + ß. &= K. [.067 * 2.0247 B. = 3.4810-4. &= k. [oista prost1419 = 947 of reaction Hence. to NO according order = 1 Similerly & = 8.5x10r = k & = 3:4810-4 = k [.0157 (10247 [10157* [.ouIB = 2 Hence overall order of reaction - 242 - 2 order 1 of reation wit ... No 1 Uz = 2 ra k. (Not SuJ2 3.4810-4=k. 1.067 [10